a constant torque of 10m•N is applied to the rim of a 8.0-kg uniform disk of radius 0.15m. what is the angular speed of the disk about an axis through its center after it rotates 2.0 revolutions from rest?
Answer #1

torque = 10 Nm


mass m = 8 kg


radius = 0.15 m

angular displacemet theta = 2 revolution = 2*2*pi rad


work done = change in KE


torque *displacemenet = 0.5*I*(w2^2-w1^2)


10*2*2*pi = 0.5*(1/2)*8*0.15^2*(w2^2-0^2)

w2 = 52.844 rad/s <<----------answer